OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (texaschic101):

if you are looking for a parallel line, the slope is going to be the same. So, lets find the slope.... put your equation in y = mx + b and m will be your slope. Can you find the slope ?

OpenStudy (texaschic101):

are you having trouble ?

OpenStudy (texaschic101):

tag me when you are ready to participate

OpenStudy (anonymous):

x + 4y = 6, 4y=-x+6, divide everything by 4 .